Suppose that a competitive firm’s marginal cost of producing output q is given

by MC(q) = 3 + 2q.

Assume that the market price of the firm’s product is $9.

a) What level of output will the firm produce?

b) What is the firm’s producer surplus? Please compute and illustrate graphically.


Suppose that the average variable cost of the firm is given by AV C(q) = 3 + q. Suppose that the firm’s fixed costs are known to be $3.

c) What are the variable cost ? Verify by checking the total cost function’s marginal cost !

d) Will the firm be earning a positive, negative, or zero profit in the short run?



1
Expert's answer
2021-12-03T08:34:09-0500

MC=3+2q

p=9

a) A perfectly competitive firm will produce where

P=MC

9=3+2q

"q^* = \frac{9-3}{2} = 3 \; units \\

p^* = \%9"

b) Imperfect of MC curve = 3 + 2(0) = 3

PS = "\frac{1}{2}" (Market price -Min. Price willing for receive) "\times" Q

"= \frac{1}{2}(9-3) \times 3 \\

= \$9"

c) Economic profit = TR -TC

"= (9 \times 3) -(3 +(3+3) \times 3) \\

= \$6"

d) It will earn positive economic profit in the short run.